Hello my question is:
Solid ammonium bromide is slowly added to 50.0 mL of a 0.0407 M silver nitrate solution. The concentration of bromide ion required to just initiate precipitation is?
1
Expert's answer
2020-11-12T07:11:45-0500

The chemical equation for the reaction is;

"NH_4Br_{(aq)}+AgNO_{3{(aq)}}\to NH_4NO_{3{(aq)}}+AgBr_{(s)}"

The ionic equation for the reaction is;"NH_4^++Br^-+Ag^++NO_3^-\to NH_4^++NO_3^-+AgBr"

"\therefore" Net ionic equation is;

"Br^-_{(aq)}+Ag^+_{(aq)}\to AgBr_{(s)}"

"K_{sp}of AgNO_3=1.8\times 10^{-18}"

"K_{sp} of AgBr=7.7\times 10^{-13}"

"K_{sp}AgNO_3=[Ag^+][NO_3^-]"

"1.8\times 10^{-18}=[Ag^+]\times 0.0407M"

"[Ag^+]" "=4.4226\times 10^{-18}M"

"K_{sp}AgBr=[Ag^+][Br^-]"

"7.7\times 10^{-13}=" "4.4226\times 10^{-18}" "\times[Br^-]"

"[Br^-]=1.741\times10^4M"

"\therefore" The concentration of "Br^-=" "1.741\times10^4M"